OpenStudy (anonymous):

find the slope of the line passing through the pair of points or state that the slope is undefined. Then indicate whetherthe line through the pair rises, falls, is horizontal, or is vertical (-4,4)and (5,3)

OpenStudy (anonymous):

slope or gradient =(Y2-Y1)/(X2-X1) =(3-4)/(5-4) = -1/1 =-1 since the slope is negative the line falls

OpenStudy (anonymous):

any points 4and3 can be taken as y1 or y2 and also the same for -4 and 5

OpenStudy (anonymous):

thank you